OpenStudy (anonymous):

sec^2 18degrees- tan^2 18degrees

OpenStudy (campbell_st):

use sin^2 + cos^2 = 1 divide by cos^2 gives tan^2 + 1 = sec^2 or sec^2 = tan^2 - 1 so its tan^2(18) - 1 - tan^2(18)